# Coupled Interfacial Phenomena Suppress Propulsion in Catalytic Janus Colloids

### v1

@@ Abstract
Introduction - Active matter is classified as a system that extracts energy from its environment to exert force and exhibit motion[1]. Examples of active matter include microscopic living organisms, robots, and motors[2], all of which are typically at low Reynolds numbers where viscous forces dominate inertia. Often complementing the active force is that of stochastic ”Brownian” fluctuations in both position and orientation that are relevant considering deterministic active motion. Various external applied or locally generated fields, including electric[3], magnetic [4], optical[5], acoustic [6], thermal[7], or chemical gradients [8], are used to actuate motion. In addition to serving as excellent experimental models for condensed matter physics, active particles are thought to have potential in diverse applications, including environmental remediation, micro-scale cargo transport[9, 10, 11], microsurgery [12], bio-sensing [13], water purification [14], oil recovery [15], and drug delivery [16]. Active particles also serve as model systems for studying biological propulsion[17]. In both fundamental studies and applied uses of active particles, interfacial phenomena involving the particle and nearby boundaries play a critical role. For instance, microorganisms often collect near interfaces (e.g. ocean vegetation or solid surfaces) making it essential to understand how boundaries influence their motion.[18].
@@
An emerging theme among studies over the past decade has been the importance of interfacial properties of either the particle itself or nearby boundary. Proximity and orientation with respect to a boundary, as well as interfacial chemistry, has been shown to significantly impact propulsion speed. For example, Janus particles near boundaries tend to move in a sliding state due to hydrodynamic interactions, gravitational torque on cap, and sedimentation of particles [21, 24, 8]. Moreover, the angle at which active particle approaches a boundary can also effect whether it moves along the boundary, is reflected back, or becomes stationary[25, 26]. Particles can slide along both bottom and top boundaries of a fluid cell, referred to as flooring and ceiling states respectively as shown in Figure 1 a [21]. Interesting recent work in this area has shown that altering the tilt angle of the platinum cap with respect to the boundary using an external electric field increases the particle velocity[27]. Boundary chemistry can also influence propulsion. For example, particles move faster near hydrophobic surfaces than hydrophilic ones, due to osmotic coupling with the boundary [28]. Note that plasma treatment is regularly used as a standard cleaning protocol for cleaning the substrate for Janus particle experiments, known to make the surface hydrophilic and negatively charged [24], which suggests that surface preparation may play a significant role in many of the existing studies.
@@
Coupling to the phenomena described above is the geometry (coverage and thickness) and composition (material) of the cap itself. Modeling of non-active systems suggest that increasing cap thickness enhances the likelihood of a quenched state (cap pointing downward), where gravitational torque from the denser platinum (22.50 g/cm³) compared to polystyrene (1.05 g/cm³) suppresses rotational diffusion[29, 30, 31, 32] . There has been considerably fewer studies that consider the effect of cap thickness in active systems [25]. Yet, the modeling of non-active systems referenced above showed that even caps of moderate thickness will significantly alter the orientational dynamics of a Janus particle. One study showed that particle velocities decrease when the cap thickness is 5 nm due to lower coverage of the catalytic material on the particle surface, thus lowering the overall reaction rate on the platinum cap[33]. Several experimental studies have used different cap thicknesses between 5 nm and 20 nm and it is reported that platinum thickness is typically non-uniform, forming a crescent-like shape [21, 34, 35, 36, 37]. Theoretical predictions agree with previous experimental observations of caps orienting towards the bottom boundary in the absence of hydrogen peroxide in solution [8]. Upon addition of peroxide, particles slide in parallel along the bottom and top surfaces, slightly tilted away from the boundary [38]. The average tilt of active particles is found to be approximately 105 ∘ as illustrated in Figure 1 d. Further, it is shown that bottom-heaviness of a platinum cap can cause active particles to propel against gravity[39]. While Wang et al. showed velocity increases with increasing platinum cap thickness from 5 nm to 20 nm, their study did not specify whether measurements were conducted near a boundary. Further, the effect of cap thickness was outside of the scope of that initial study, whereas it instead explored difference arising from the surface contour of platinum in the cap (e.g. smooth vs. nanoparticle platinum) [40]. A more recent study has shown the velocities of a Janus particle near an indium-tin oxide (ITO) coated glass slide and in 5 v/v. % peroxide will increase to as much as 5 m/s [38]. One clever strategy in their work was to apply an intermittent A/C electric field, which was used to help control particle orientation but it was unclear what impact application of en electric field would have on equilibrium or transient charge on the particle or ITO surfaces.
@@
Flooring and ceiling dynamics as a function of cap thickness - Janus particles suspended in peroxide solution undergo a combination of stochastic and deterministic motion that, overtime, results in accumulation of particles at either the top (i.e. ceiling) or bottom (i.e. flooring) boundaries, as illustrated in Figure 1. While theoretical studies have predicted the bifurcation of populations into flooring and ceiling states, few experimental studies have simultaneously observed both. Herein, we present measurements to gauge the importance of tracking both populations. Figure 2 shows measurements of velocity and maximum displacement for both flooring and ceiling populations as a function of cap thickness. Interestingly, although there is a dependence on cap thickness (as discussed later), there is no measurable difference in velocity or displacement between the flooring and ceiling populations. For all experimental systems, from cap thicknesses of 3 nm to 35 nm, flooring and ceiling populations propel with statistically similar velocities and experience similar displacements. There remains lack of consensus in the community, in that Das et al. have argued that ceiling and flooring are similar, whereas Wang et al. presented data showing that the flooring velocity is higher than the ceiling velocity due to lower tilt on the floor[21, 38].
@@
Conclusion - In summary, this study provides a systematic experimental investigation of Janus particle dynamics as a function of platinum cap thickness, peroxide concentration, and substrate treatment. While theoretical studies have long predicted bifurcation into ceiling and flooring states, our results confirm that both populations exhibit statistically similar velocities and displacements across a wide range of cap thicknesses, thereby underscoring the importance of tracking both states simultaneously. We demonstrate that propulsion strongly depends on platinum cap thickness and peroxide concentration, with thinner caps () limited by incomplete catalytic coverage, intermediate caps () yielding peak velocities and displacements, and thicker caps () showing reduced activity due to gravitational torque and reorientation effects. Plasma cleaning of substrates revealed an unexpected global suppression of particle velocities at both boundaries, consistent with charge diffusion from the treated surface into solution, thereby establishing a new link between substrate conditioning and active particle propulsion. Collectively, these findings not only extend experimental validation of ceiling and flooring states but also uncover how catalytic layer thickness, chemical environment, and boundary conditions couple to determine active particle trajectories. This work provides critical insight into how Janus particle motion can be tuned and controlled, with implications for the rational design of active colloids in applications ranging from targeted delivery to boundary-mediated assembly.

### v2

@@ Abstract
Introduction - Active matter is classified as a system that extracts energy from its environment to exert force and exhibit motion[1]. Examples of active matter include microscopic living organisms, robots, and motors[2], all of which are typically at low Reynolds numbers where viscous forces dominate inertia. Often complementing the active force is that of stochastic ”Brownian” fluctuations in both position and orientation that are relevant considering deterministic active motion. Various externally applied or locally generated fields, including electric[3], magnetic [4], optical[5], acoustic [6], thermal[7], or chemical gradients [8], are used to actuate motion. In addition to serving as excellent experimental models for condensed matter physics, active particles are thought to have potential in diverse applications, including environmental remediation, micro-scale cargo transport[9, 10, 11], microsurgery [12], bio-sensing [13], water purification [14], oil recovery [15], and drug delivery [16]. Active particles also serve as model systems for studying biological propulsion[17]. In both fundamental studies and applied uses of active particles, interactions involving the particle and nearby boundaries play a critical role. For instance, microorganisms often collect near interfaces (e.g. ocean vegetation or solid surfaces) making it essential to understand how boundaries influence their motion.[18].
@@
An emerging theme among studies over the past decade has been the importance of interfacial properties of either the particle itself or nearby boundary. Proximity and orientation with respect to a boundary, as well as interfacial chemistry, have been shown to significantly impact propulsion speed. For example, Janus particles near boundaries tend to move in a sliding state due to hydrodynamic interactions, gravitational torque on cap, and sedimentation [21, 24, 8]. Moreover, the angle at which an active particle approaches a boundary can effect whether it moves along the boundary, is reflected back, or becomes stationary[25, 26]. Particles can slide along both bottom and top boundaries of a fluid cell, referred to as flooring and ceiling states respectively, as shown in Figure 1 a [21]. Interesting recent work in this area has shown that altering the tilt angle of the platinum cap with respect to the boundary using an external electric field increases the particle velocity[27]. Boundary chemistry can also influence propulsion. For example, particles move faster near hydrophobic surfaces than hydrophilic ones, due to osmotic coupling with the boundary [28]. Note that plasma treatment is regularly used as a standard cleaning protocol for cleaning the substrate for Janus particle experiments, known to make the surface hydrophilic and negatively charged [24], which suggests that surface preparation may play a significant role in many of the existing studies.
@@
Complementing the boundary mediated phenomena described above is the geometry (coverage and thickness) and composition (material) of the cap itself. Modeling of non-active systems suggest that increasing cap thickness enhances the likelihood of a quenched state (cap pointing downward), where gravitational torque from the denser platinum (22.50 g/cm³) compared to polystyrene (1.05 g/cm³) suppresses rotational diffusion[29, 30, 31, 32] . There has been considerably fewer studies that consider the effect of cap thickness in active systems [25]. Yet, the modeling of non-active systems referenced above showed that even caps of moderate thickness will significantly alter the orientational dynamics of a Janus particle. One study showed that particle velocities decrease when the cap thickness is 5 nm due to lower coverage of the catalytic material on the particle surface, thus lowering the overall reaction rate on the platinum cap[33]. Several experimental studies have used different cap thicknesses between 5 nm and 20 nm and it is reported that platinum thickness is typically non-uniform, forming a crescent-like shape [21, 34, 35, 36, 37]. Theoretical predictions agree with previous experimental observations of caps orienting towards the bottom boundary in the absence of hydrogen peroxide in solution [8]. Upon addition of peroxide, particles slide in parallel along the bottom and top surfaces, slightly tilted away from the boundary [38]. The average tilt of active particles is found to be approximately 105 ∘ as illustrated in Figure 1 d. Further, it is shown that bottom-heaviness of a platinum cap can cause active particles to propel against gravity[39]. While Wang et al. showed velocity increases with increasing platinum cap thickness from 5 nm to 20 nm, their study did not specify whether measurements were conducted near a boundary. Further, the effect of cap thickness was outside of the scope of that initial study, whereas it instead explored differences arising from the surface contour of platinum in the cap (e.g. smooth vs. a nanoparticle platinum film) [40]. A more recent study has shown the velocities of a Janus particle near an indium-tin oxide (ITO) coated glass slide and in 5 v/v. % peroxide will increase to as much as 5 m/s [38]. One of the clever strategies in their work was to apply an intermittent A/C electric field, which was used to help control particle orientation. However, it was unclear what impact application of en electric field would have on equilibrium or transient charge on the particle or ITO surfaces.
@@
Flooring and ceiling dynamics as a function of cap thickness - Janus particles suspended in peroxide solution undergo a combination of stochastic and deterministic motion that, overtime, results in accumulation of particles at either the top (i.e. ceiling) or bottom (i.e. flooring) boundaries, as illustrated in Figure 1. While theoretical studies have predicted the bifurcation of populations into flooring and ceiling states, few experimental studies have simultaneously observed both. Herein, we present measurements to gauge the importance of tracking both populations. Figure 2 shows measurements of velocity and maximum displacement for both flooring and ceiling populations as a function of cap thickness. Interestingly, although there is a dependence on cap thickness (as discussed later), there is no measurable difference in velocity or displacement between the flooring and ceiling populations. For all experimental systems, from cap thicknesses of 3 nm to 35 nm, flooring and ceiling populations propel with statistically similar velocities and experience similar displacements. Note there remains lack of consensus in the community, in that Das et al. have argued that ceiling and flooring are similar, whereas Wang et al. presented data showing that the flooring velocity is higher than the ceiling velocity due to lower tilt on the floor[21, 38].
@@
Conclusion - In summary, this study provides a systematic experimental investigation of Janus particle dynamics as a function of platinum cap thickness, peroxide concentration, and substrate treatment. While theoretical studies have long predicted bifurcation into ceiling and flooring states, our results confirm that both populations exhibit statistically similar velocities and displacements across a wide range of cap thicknesses, thereby underscoring the importance of tracking both states simultaneously. We demonstrate that propulsion strongly depends on platinum cap thickness and peroxide concentration, with thinner caps () limited by incomplete catalytic coverage, intermediate caps () yielding peak velocities and displacements, and thicker caps () showing reduced activity due to gravitational torque and reorientation effects. Plasma cleaning of substrates revealed an unexpected global suppression of particle velocities at both boundaries, consistent with charge diffusion from the treated surface into solution, thereby establishing a new link between substrate conditioning and active particle propulsion. Collectively, these findings not only extend experimental validation of ceiling and flooring states but also uncover how catalytic layer thickness, chemical environment, and boundary conditions couple to determine active particle trajectories. This work provides critical insight into how Janus particle motion can be tuned and controlled, with implications for the rational design of active colloids.